# Branch structure and nonextensive thermodynamics of Kalb-Ramond-ModMax black holes: observational signatures

### v1

@@ VIII Conclusions
Optical properties provide an independent and complementary window into the branch structure. Using the Gauss-Bonnet approach to weak gravitational lensing, we found that the ordinary branch generally produces reduced deflection angles compared to the Schwarzschild case, reflecting the partially repulsive role of the electromagnetic sector. In contrast, the phantom branch yields systematically enhanced deflection, since the sign-reversed electromagnetic contribution reinforces gravitational focusing. This distinction persists, in a modified form, when photon propagation is studied in plasma environments. In homogeneous plasma, branch-dependent shifts of the photon-sphere radius remain visible, while in strongly stratified plasma distributions the optical structure becomes dominated by the Lorentz-symmetry-breaking parameter $\ell$, with ModMax effects largely suppressed. This hierarchy reveals the importance of accurate plasma modeling when attempting to extract fundamental physics from black-hole shadow observations.

### v2

@@ VIII Conclusions
Optical properties provide an independent and complementary window into the branch structure. Since the KR-ModMax metric satisfies $f(r)\to 1/(1-\ell)\neq 1$ at spatial infinity, the standard Gibbons-Werner formulation of the Gauss-Bonnet theorem is inapplicable. We therefore employed the Ono-Ishihara-Asada extension, which introduces a topological boundary correction $\pi(\sqrt{1-\ell}-1)<0$ that universally reduces light bending relative to the Schwarzschild baseline. The effective gravitational mass $M(1-\ell)$ in the leading term further suppresses deflection, a qualitative departure from the Letelier cloud-of-strings spacetime, where the analogous topological term is positive and enhances light bending. This sign reversal reflects the distinct origin of non-flatness: the KR condensate modifies the lapse rather than the angular metric, producing an effective angular surplus in the optical geometry. The branch dichotomy $\zeta=\pm 1$ persists through the charge-dependent second-order terms, with the phantom branch yielding larger deflection angles due to the constructive alignment of mass and charge contributions. This distinction persists, in a modified form, when photon propagation is studied in plasma environments. In homogeneous plasma, branch-dependent shifts of the photon-sphere radius remain visible, while in strongly stratified plasma distributions the optical structure becomes dominated by the Lorentz-symmetry-breaking parameter $\ell$, with ModMax effects largely suppressed. This hierarchy reveals the importance of accurate plasma modeling when attempting to extract fundamental physics from black-hole shadow observations.